OpenStudy (anonymous):

A 73.0-kg person stands on a scale in an elevator. What is the apparent weight when the elevator isaccelerating upward with an acceleration of 1.00 m/s2,

OpenStudy (anonymous):

We know g is 9.8 m/s, unless you are on another planet.. In this case, safe to say we are on earth. If you are going upwards, you can feel more pressure in your feet, tell tale sign that there is an increase downward force.. F = m*a where a = g + added g from elevator. effectively putting you at a = 10.8 m/s^2 in the downward direction. Take you're mass and multiply it by the acceleration and you have the force in newtons: If you can guess the equation...... F = 76kg*10.8 = 820.8 Newtons... IMPORTANT.... They say what is the apparent weight which is a tricky thing.. So take that 820.8 divid it by 9.8, you get 83.8 kg.. That means that the person feels 83.8-76 = 7.8 kg heavier... Becareful on what the question is asking for

OpenStudy (anonymous):

oops.. I wrote 76 kg for the person's weight.. go back and change that in the calculation.

OpenStudy (anonymous):

so is the answer 83.8 or 7.8?

OpenStudy (anonymous):

If you are going upward you should appear lighter.

OpenStudy (anonymous):

Since you missed my second comment.. I will recalcuate it for you.. The person's weight would feel as though it's 80.4kg